OpenStudy (anonymous):

if f(x) = −5x + 1, find f(−3).

OpenStudy (anonymous):

-5(-3)+1=16

OpenStudy (anonymous):

-5(-3)+1 negatives cancel: 5(3)+1 15+1 16

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

f(-3) = -5(-3) +1 f(-3) = 15 + 1 f(-3) = 16